# Pre-training Differentially Private Models with Limited Public Data

### v1

@@ Abstract
The superior performance of large foundation models relies on the use of massive amounts of high-quality data, which often contain sensitive, private and copyrighted material that requires formal protection. While differential privacy (DP) is a prominent method to gauge the degree of security provided to the models, its application is commonly limited to the model fine-tuning stage, due to the performance degradation when applying DP during the pre-training stage. Consequently, DP is yet not capable of protecting a substantial portion of the data used during the initial pre-training process.
In this work, we first provide a theoretical understanding of the efficacy of DP training by analyzing the per-iteration loss improvement. We make a key observation that DP optimizers’ performance degradation can be significantly mitigated by the use of limited public data, which leads to a novel DP continual pre-training strategy. Empirically, using only 10% of public data, our strategy can achieve DP accuracy of 41.5% on ImageNet-21k (with $\epsilon=8$), as well as non-DP accuracy of 55.7% and and 60.0% on downstream tasks Places365 and iNaturalist-2021, respectively, on par with state-of-the-art standard pre-training and substantially outperforming existing DP pre-trained models.
@@ 1 Introduction
Pre-trained large models have been the backbone of computer vision and natural language processing. Finetuning or zero/few-shots learning (including in-context learning) based on these models can achieve superior performance. In particular, differentially private (DP) fine-tuning, such as full-parameter training, LoRA, Adapter, BiTFiT, and linear probing  [88, 48, 11, 12, 57], has shown to be almost as accurate as the standard non-DP fine-tuning on GPT [67, 6], ViT [26], and ResNet [38] models, while protecting the privacy of fine-tuning data privacy. To be more specific, DP language models (GPT2/BERT/RNN) and vision models are highly effective in defending against canary insertion attacks [65, 39, 41] and membership inference attacks [19, 68, 31, 15]; DP fine-tuned GPT2 also reduces the personally identifiable information leakage by $5-10$ times compared with its Non-DP counterpart[52].
These DP fine-tuned models all follow a two-stage training procedure, in which the first stage trains a pre-trained model on large public datasets (without privacy protection), and the second stage fine-tunes on relatively small private datasets. However, a growing concern has been raised against the pre-training on the vast amount of web-collected data [17, 16, 77, 40, 61, 66]. The pre-trained models could memorize and re-generate the sensitive information in the training data, which include copyright content in books and codes, images of faces and nudity, and other personally identifiable information such as address and phone numbers, even if the data have been pre-processed by some content filters. While it is common to use close-source or proprietary datasets instead and not to release the models trained on them, this approach renders the models not reproducible and still violates data privacy in a more implicit way.
Consequently, because of the uncertainty in seemingly safe-to-use public datasets, it is important to apply DP to the pre-training phase, which is computationally feasible, since DP optimization can be as efficient as the standard Non-DP training [13, 37]. However, DP optimization without any public data suffers from slow convergence, and suboptimal performance: even on CIFAR10, the accuracy drops from $>95\%$ to $<70\%$ at $\epsilon=8$ [62]; on GPT2, the non-DP BLEU score degrades from 65.73 (quality often better than human) to 15.457 (hard to get the gist) at $\epsilon=3$ [48]. In short, pre-training and fine-tuning differ significantly in the amount of data and computation resources used, as well as in optimization setups. We summarize the difference between DP pre-training and DP fine-tuning in Table 1.
@@
Public data size used small large
@@
Trainable parameters 100% 0.1-100%
@@ 1.1 Contributions
Our main contributions are listed below.
@@
(2) Based on the framework, we can analyze the effect that DP mechanisms, especially the DP noise, have on pre-training and fine-tuning stages, leading to some theoretical justifications about why pre-training is more vulnerable to DP noise; see Figure 1 for some empirical comparisons, and Section 3 for some in-depth discussion.
(3) Our analysis suggests that the deceleration due to DP mechanisms can be mitigated by using a certain amount of public data. We then propose a DP continual pre-training strategy that continues the pre-training privately from a non-DP initialization. Our strategy is easily implementable, automatic, and effective, which obtains accuracy on upstream and downstream tasks that is on par with its non-DP variants, while significantly outperforming state-of-the-art DP pretraining algorithms (see Figure 2). We will open-source our DP pre-trained vision models.
@@
Figure 1: Comparison among the convergence of standard SGD, clipped SGD, noisy SGD, and DP-SGD in different tasks and training stages.
Figure 2: Summary of results in Section 5. First three figures compare the downstream and few-shot performance and the data efficiency (circle’s radius proportional to pre-training data size) of the DP pre-trained vision models; the last figure shows the performance of DP pre-trained models defending against privacy attacks (lower is more private).
@@ 1.3 Notions & settings
We study DP optimization under fixed budgets for computation (with a fixed number of samples), privacy (with a fixed DP guarantee), and limited availability of public data.
@@ 1.4 Computation budget
As a consequence, the batch size $B$ has nearly zero influence on the total training time, focusing its effect only on the convergence speed.
@@ 2 Understanding DP training through the lens of Hessian
In this section, we derive the impact of different components in DP training (including the per-sample clipping, the noising, and hyper-parameter choices), by characterizing the per-iteration per-sample loss improvement through the lens of Hessian.
To begin with, we consider the task of optimizing the generalization loss (equivalently the expected risk) for any differentiable loss function, e.g. the cross entropy or adversarial loss [32]:
@@
Consider the general DP optimizers, including SGD and Adam [45], which update the parameters with the privatized gradient,
@@
Figure 4 illustrates the values of $C_{i}$ under different clipping functions. Note that in (2), we employ a crucial approximation $c\approx\mathbb{E}[C_{i}]$ so as to unify the formula of DP and non-DP SGD in Remark 2.2. This approximation only holds when the directions of vectors $\sum_{i}C_{i}{\mathbf{g}}_{i}$ and $\sum_{i}{\mathbf{g}}_{i}$ are very close, i.e., there is little per-sample clipping bias. Such approximation is validated in Figure 1, where we observe from the ‘SGD’ and ‘SGD+clipping’ curves that the convergence (without noising) is not much influenced by the bias of per-sample gradient clipping.
@@
Setting $c=1$ and $\sigma=0$, the gradient (2) reduces to the standard minibatch gradient. Hence, the difference between SGD and DP-SGD is characterized by $(\sigma,c)$.
Per-iteration improvement of DP-SGD:
Next, we characterize and analyze the per-iteration improvement of DP-SGD through the lens of Hessian, under different choices of hyperparameters and clipping functions:
${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta{\mathbf{g}}_{t}.$ (4)
The extension of the analysis to more general optimizers (e.g., DP-Adam) is given in Section 3.4. We are interested in minimizing the second-order Taylor expansion of $L({\mathbf{w}}-\eta{\mathbf{g}})$, which is sufficiently accurate since parameter updates are often quite small [56]. The loss improvement in one iteration can be approximated as:
@@
Taking the expectation, we obtain the expected per-iteration loss improvement (derived in Section A.1):
@@
By applying Assumption 2.1, we have
$\mathbb{E}[{\mathbf{g}}]=c{\mathbf{G}},\quad\textup{Cov}({\mathbf{g}})=c^{2}\mathbf{\Sigma}/B+\sigma^{2}/B^{2}.$
Substitute to (5), we obtain a quadratic function of $\eta$:
@@
By optimizing the learning rate $\eta$ 222 The optimal learning rate of DP-SGD cannot be used in practice because (i) the oracle ${\mathbf{G}}$ and ${\mathbf{H}}$ are unknown; (ii) it is data-dependent and hence violates the DP guarantee. However, in our analysis such an optimal learning rate helps us understand the best possible per-iteration performance. , the per-sample, per-iteration improvement, $\max_{\eta}\Delta L_{\textup{priv}}/B$, simplifies to
$\Delta^{\star}_{\textup{priv}}(B):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})}.$ (7)
Given that the total number of processed samples is fixed at $S=BT$, (7) can be used as a metric to evaluate the data efficiency of different training settings, e.g. DP and non-DP training with different $T,B$ and $(\epsilon,\delta)$.
Per-iteration improvement of vanilla SGD: According to Remark 2.2, we can analyze the loss improvement of standard SGD (which uses full public data) as a sub-case of DP-SGD by substituting $c=1,\sigma=0$ into (6). Additionally, we denote the batch size used for SGD and DP-SGD as $B_{\text{non-DP}}$ and $B_{\text{DP}}$, respectively:
$\Delta L_{\textup{pub}}$ $:=\eta{\mathbf{G}}^{\top}{\mathbf{G}}-\frac{\eta^{2}}{2}\left(\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B_{\text{non-DP}}}+{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\right).$ (8)
Applying the optimal learning rate $\eta$, the per-sample per-iteration improvement $\max_{\eta}\Delta L_{\textup{pub}}/B_{\text{non-DP}}$ is
$\Delta^{\star}_{\textup{pub}}(B):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}$ (9)
@@
From (7), it is clear that smaller $\sigma$ (from the advances in privacy accounting theory) and larger $c$ (hence larger $C_{i}$) can help DP training. Under the same sensitivity bound, AUTO clipping gives the largest $C_{i}$ among all clipping functions (see Figure 4), therefore is more preferred to use in practice.
Implication 2.4 (Moderately large batch size helps DP training).
There exists an optimal batch size $B^{\star}_{\text{DP}}:=\text{argmax}_{B}\Delta^{\star}_{\textup{priv}}(B)\approx\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}$, as visualized in Figure 5. Compared to previous DP literatures, our derivation of $B^{\star}_{\text{DP}}$ also supports the use of large batch size, but further indicates the data inefficiency if $B_{\text{DP}}$ is too large, a case that is often overlooked.
@@ 3 Impact of per-sample clipping and noising
Comparing DP-SGD (7) to SGD (9), we can attribute the slow convergence of DP optimization to the term $\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})$ in (7), which is not present in the standard training. We referred to such a term as the “decelerator”:
@@
Next, we discuss how the decelerator impacts the non-DP training, DP pre-training, and DP fine-tuning stages.
@@ 3.2 DP pre-training can be vulnerable to noise
When $\sigma\neq 0$, the decelerator is non-zero. Therefore, DP training is slowed down by the noise; see Figure 1 that SGD with noise (yellow and red curves) has worse performance than SGD without noise (black and blue curves). Furthermore, in pre-training, the decelerator is relatively large in the denominator of (7), i.e., $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})\leq\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{Bc^{2}}$ unless $B$ is large which is computationally inefficient (see Figure 5, upper left plot), and therefore, the slowdown can be significant.
Note that the deceleration issue cannot be resolved by increasing $B$. Although increasing $B$ improves the relative (to non-DP) speed of DP convergence, i.e., the decelerator decreases, it hurts the absolute speed since $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}$ increases, and thus the loss improvement (7) also worsens (c.f. Figure 5 upper row). Therefore, to design an efficient DP pre-training strategy, we must keep $B$ moderate and reduce the decelerator simultaneously.
Figure 5: Illustration of different terms (7) for DP pre-training (upper) and DP fine-tuning (lower). Left sub-plots depict the denominator in (7).

### v2

@@ Abstract
The superior performance of large foundation models relies on the use of massive amounts of high-quality data, which often contain sensitive, private and copyrighted material that requires formal protection. While differential privacy (DP) is a prominent method to gauge the degree of security provided to the models, its application is commonly limited to the model fine-tuning stage, due to the performance degradation when DP is applied during the pre-training stage. Consequently, DP is yet not capable of protecting a substantial portion of the data used during the initial pre-training process. In this work, we provide a theoretical understanding of the efficacy of DP training by analyzing the per-iteration loss improvement, through the lens of Hessian matrix for large neural networks. We make a key observation that DP optimizers’ performance degradation can be significantly mitigated by the use of limited public data, which leads to a novel DP continual pre-training strategy. Empirically, using only 10% of public data, our strategy can achieve DP accuracy of 41.5% on ImageNet-21k (with $\epsilon=8$), as well as non-DP accuracy of 55.7% and 60.0% on downstream tasks Places365 and iNaturalist-2021, respectively, on par with state-of-the-art standard pre-training and substantially outperforming existing DP pre-trained models. Our DP pre-trained models are released in fastDP library (https://github.com/awslabs/fast-differential-privacy/releases/tag/v2.1).
@@ 1 Introduction
Large pre-trained models have been the backbone of computer vision and natural language processing. Finetuning or zero/few-shots learning (including in-context learning) based on these models can achieve superior performance. In particular, differentially private (DP) fine-tuning, such as full-parameter training, LoRA, Adapter, BiTFiT, and linear probing [89, 49, 14, 12, 58, 24], has shown to be almost as accurate as the standard non-DP fine-tuning on GPT [68, 6], ViT [27], and ResNet [39] models, while protecting the privacy of fine-tuning data. To be more specific, DP language models (GPT/BERT/RNN) and vision models are highly effective in defending against canary insertion attacks [66, 40, 42] and membership inference attacks [20, 69, 32, 16]; DP fine-tuned GPT2 also reduces the personally identifiable information leakage by $5\sim 10$ times compared with its non-DP counterpart [53].
These DP fine-tuned models all follow a two-stage training procedure, in which the first stage trains a model on large public datasets from scratch (without privacy protection), and the second stage fine-tunes on relatively small private datasets. However, a growing concern has been raised against the pre-training on the vast amount of web-collected data [18, 17, 79, 41, 62, 67]. The pre-trained models could memorize and re-generate the sensitive information in the training data, which include copyright content in books and codes, images of faces and nudity, and other personally identifiable information such as address and phone numbers, even if the data have been pre-processed by some content filters. While it is common to use close-source or proprietary datasets (e.g. JFT [75, 24, 58]) instead and to not release the models trained on these datasets, this approach renders the models not reproducible and may still violate the data privacy in a more implicit way.
Consequently, because of the uncertainty in seemingly safe-to-use public datasets, it is important to apply DP to the pre-training phase, which is computationally feasible, since DP optimization can be as efficient as the standard non-DP training [13, 38]. However, DP optimization without any public data suffers from slow convergence, and suboptimal performance: even on CIFAR10, the accuracy drops from $>95\%$ to $<70\%$ at $\epsilon=8$ [63]; on GPT2, the non-DP BLEU score degrades from 65.73 (quality often better than human) to 15.457 (hard to get the gist) at $\epsilon=3$ [49]. In short, pre-training and fine-tuning differ significantly in the amount of data and computation resources used, as well as in optimization setups. We summarize the difference in Table 1.
@@
(amount of compute)
Trainable model parameters 100% $0.1\sim 100$%
Major cause of DP noising DP clipping
performance degradation
@@ 1.1 Contributions
Our main contributions are listed below. We emphasize that we have taken the privacy accounting and the convergence (i.e. the dependence of loss on $\sigma(B,T,\epsilon),B$) jointly into consideration.
@@
(2) Based on the framework, we can analyze the effect that DP mechanisms, especially the DP noise but less so the DP clipping, have on pre-training and fine-tuning stages, leading to some theoretical justifications about why pre-training is more vulnerable to DP noise; see Figure 1 for some empirical comparisons, and Section 3 for some in-depth discussion.
(3) Our analysis suggests that the deceleration due to DP mechanisms can be mitigated by using a certain amount of public data. We then propose a DP continual pre-training strategy that continues the pre-training privately from a non-DP initialization. Our strategy is easily implementable, automatic, and effective, which obtains accuracy on upstream and downstream tasks that is on par with its non-DP variants, while significantly outperforming state-of-the-art DP pretraining algorithms (see Figure 2).
@@
Figure 1: Comparison among the convergence of standard SGD, clipped SGD without noise, noisy SGD without clipping, and DP-SGD in different tasks and training stages.
Figure 2: Summary of results in Section 5. First three figures compare the downstream and few-shot performance and the data efficiency (circle’s radius proportional to pre-training data size) of the DP pre-trained models; the last figure shows the performance of DP pre-trained models defending against privacy attacks (lower is stronger in defense).
@@ 1.3 Notions & settings
We study DP optimization under a fixed computation budget (with a fixed number of samples), a fixed privacy budget (with a fixed DP guarantee), and limited availability of public data.
@@ 1.4 Computation budget
As a consequence, the batch size $B$ has nearly zero influence on the total training time, leaving its effect only on the convergence speed.
@@ 2 Understanding DP training through the lens of Hessian
In this section, we derive the impact of different components in DP training, including the per-sample clipping, the noising, and hyper-parameter choices. To do so, we characterize the per-iteration per-sample loss improvement through the lens of Hessian.
We aim to optimize the generalization loss (equivalently the expected risk) for any differentiable loss function, e.g. the cross entropy, mean squared error, or adversarial loss [33]:
@@
Consider the general DP optimizers, such as SGD and Adam [46], which update the parameters with the privatized gradient,
@@
Figure 4 illustrates the values of $C_{i}$ under different clipping functions. Note that in (2), we employ a crucial approximation $c\approx\mathbb{E}[C_{i}]$ so as to unify the formula of DP and non-DP SGD in Remark 2.2. This approximation only holds when the directions of vectors $\sum_{i}C_{i}{\mathbf{g}}_{i}$ and $\sum_{i}{\mathbf{g}}_{i}$ are very close, i.e., there is little per-sample clipping bias. Such approximation is empirically validated in Figure 1, where we observe from the ‘SGD’ and ‘SGD+clipping’ curves that the convergence (without noising) is not much influenced by the bias of per-sample gradient clipping.
@@
Setting $c=1$ and $\sigma=0$, the gradient (2) reduces to the standard mini-batch gradient. Hence, the difference between SGD and DP-SGD is characterized by $(\sigma,c)$.
2.1 Per-iteration improvement of DP-SGD
Next, we characterize and analyze the per-iteration improvement of DP-SGD through the lens of Hessian, under different choices of hyperparameters and clipping functions: ${\mathbf{w}}_{t+1}={\mathbf{w}}_{t}-\eta{\mathbf{g}}_{t}$. The extension of the analysis to more general optimizers (e.g., DP-Adam) is given in Section 3.4. We are interested in minimizing the second-order Taylor approximation of $L({\mathbf{w}}-\eta{\mathbf{g}})$, which is sufficiently accurate since parameter updates are often quite small [57]. The loss improvement in one iteration can be approximated as:
@@ 2.1 Per-iteration improvement of DP-SGD
Taking the expectation of the right-hand side, we obtain the expected per-iteration loss improvement (derived in Section A.1):
@@
By applying Assumption 2.1 and (2), we have $\mathbb{E}[{\mathbf{g}}]=c{\mathbf{G}},\quad\textup{Cov}({\mathbf{g}})=c^{2}\mathbf{\Sigma}/B+\sigma^{2}/B^{2}.$ Substitute to (4), we obtain a quadratic function of $\eta$:
@@
We denote the batch size used for SGD and DP-SGD as $B_{\text{non-DP}}$ and $B_{\text{DP}}$, respectively. Then by optimizing the learning rate $\eta$ 222 The optimal learning rate of DP-SGD cannot be used in practice because (i) the oracle ${\mathbf{G}}$ and ${\mathbf{H}}$ are unknown; (ii) it is data-dependent and hence violates the DP guarantee. Such an optimal learning rate is only used to help us understand the best possible per-iteration performance. , the per-sample and per-iteration improvement simplifies to
$\max_{\eta}\Delta L_{\textup{priv}}/B_{\text{DP}}=\Delta L^{\star}_{\textup{priv}}(B_{\text{DP}}):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B_{\text{DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})+\sigma^{2}\textup{tr}({\mathbf{H}})/(B_{\text{DP}}c^{2})}.$ (6)
Given that the total number of processed samples is fixed at $S=BT$, (6) can be used as a metric to evaluate the data efficiency of DP and non-DP training with different $T,B$ and $(\epsilon,\delta)$.
2.2 Per-iteration improvement of vanilla SGD
We can analyze the loss improvement of standard SGD (which uses the full public data) as a sub-case of DP-SGD by substituting $c=1,\sigma=0$ into (5), according to Remark 2.2:
$\Delta L_{\textup{pub}}$ $:=\eta{\mathbf{G}}^{\top}{\mathbf{G}}-\frac{\eta^{2}}{2}\left(\frac{\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}{B_{\text{non-DP}}}+{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}\right),\Delta L^{\star}_{\textup{pub}}(B):=\frac{1}{2}\frac{|{\mathbf{G}}|^{4}}{B_{\text{non-DP}}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})}$ (7)
We visualize (6), (7), and their individual terms in Figure 5.
(a) DP pre-training
(b) DP fine-tuning
Figure 5: Illustration of different terms in (6) and (7). Left sub-plots depict the denominators in (6) and (7). Right sub-plots depict the whole terms and optimal batch sizes.
@@ 2.2 Per-iteration improvement of vanilla SGD
From (6), it is clear that smaller $\sigma$ and larger $c$ (hence larger $C_{i}$) can help DP training. To reduce $\sigma$, we refer to Section 3.1 for a discussion of methods. For the clipping, under the same sensitivity bound, AUTO clipping [14] gives the largest $C_{i}$ among all clipping functions (see Figure 4), and therefore is more preferred to use in practice.
Implication 2.4 (Batch size should be large, but not too large).
As visualized by the red solid curves in Figure 5, there exists an optimal batch size (marked in red dashed vertical lines)
$B^{\star}_{\text{DP}}:=\text{argmax}_{B}\Delta L^{\star}_{\textup{priv}}(B)\approx\sqrt{\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{c^{2}{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}}}.$
Compared to previous DP literature, which encourages the batch size to be as a large as possible, our derivation of $B^{\star}_{\text{DP}}$ indicates a sweet pot: while we also support the use of large batch size, we highlight the data inefficiency if $B_{\text{DP}}$ is too large, a case that is often overlooked.
@@ 3 Impact of per-sample clipping and noising
Comparing DP-SGD to SGD, we can attribute the slow convergence of DP optimization to the term $\sigma^{2}\textup{tr}({\mathbf{H}})/(Bc^{2})$ in (6), which is not present in the standard training (7). We refer to such a term as
@@
Next, we discuss how the decelerator impacts the non-DP training, DP pre-training and fine-tuning.
@@ 3.1 No noise, (almost) no deceleration
Given that DP noise is critical to the convergence speed, we highlight some techniques to reduce $\sigma$ under the same budget of $(\epsilon,\delta)$: the advances in privacy accounting theory can justify smaller noise; algorithms such as LoRA, low-pass and Kalman filters [93, 92] can reduce the effective noise.
@@ 3.2 DP pre-training can be vulnerable to noise
When $\sigma\neq 0$, the decelerator is non-zero. Therefore, DP training is slowed down by the noise; in Figure 1, SGD with noise (yellow and red curves) has worse performance than SGD without noise (black and blue curves). Furthermore, in pre-training, the decelerator is relatively large in the denominator of (6), i.e., $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}+\textup{tr}({\mathbf{H}}\mathbf{\Sigma})\leq\frac{\sigma^{2}\textup{tr}({\mathbf{H}})}{Bc^{2}}$ when $B\leq B^{*}_{\text{DP}}$ (see left sub-plot of Figure 5(a)), and therefore the slowdown can be significant.
Note that the deceleration issue cannot be resolved by increasing $B$. Although increasing $B$ improves the relative speed of DP convergence in comparison to non-DP, i.e., the decelerator decreases, it hurts the absolute speed since $B{\mathbf{G}}^{\top}{\mathbf{H}}{\mathbf{G}}$ increases, and thus the loss improvement (6) also worsens (c.f. right sub-plot of Figure 5(a)). Therefore, to design an efficient DP pre-training strategy, we must keep $B$ moderate and reduce the decelerator simultaneously.